UW CHEMISTRY NMR

Announcments & Equipment Logs

December, 2015

DRX499 UPDATE – TXO PROBE INSTALLED

To all users expecting a probe swap on the drx499, the TXO 1H {13C, 19F} probe is currently installed.  

Posted in Hardware | No Comments »

AV300 ALERT! – PROBE SWAP NOW

Due to the frequently recurring 1H wobb issue, which has occurred again today, the BBI probe is being pulled and sent for repair. The new probe will be the TXD {1H, 19F, 13C} probe typically seen in the av301. Thank you for your patience. Please contact the NMR Facility Managers with any questions or concerns.

Posted in Hardware, Spectrometer | No Comments »

AV300 – UPDATE ! – SYSTEM USABLE AGAIN

We could fix the tuning problem of 1H in house.   But, I cannot emphasize enough, the following :  BE EXTREMELY GENTLE WITH THE TUNE AND MATCH RODS.  It is only a question of time before the problem recurs.  Plans are in the pipeline to provide a backup probehead for AV300 and service this probehead. […]

Posted in Hardware, Spectrometer | No Comments »

ALERT ! AV300 – 1H CHANNEL OF PROBEHEAD NON-FUNCTIONAL

The 1H channel doesn’t tune as usual. We will try to provide a workaround as soon as we can.  Please stay tuned ! NMR Facility Manager 11/2/15

Posted in Hardware, Spectrometer | No Comments »

Item Categories